OpenStudy (anonymous):

Find the cube root of 64. 2 4 8 32

OpenStudy (anonymous):

4

sam (.sam.):

\[\huge \sqrt[3]{64}=64^{\frac{1}{3}}=(4^3)^{\frac{1}{3}}\] \[\huge (4^\cancel3)^{\frac{1}{\cancel3}}=4\]